i am thinking of a 5 digit number the first digit is 2 more than the last digit but less than the second digit the second digit is the third multiple of 3 while the fourth digit is the second multiple of 3 the third digit is the difference between the second and the fourth digits what number i an thinking of?
Answer by ankor@dixie-net.com(22740)

Five digits: a,b,c,d,e
:
Write an equation for each statement
:
the first digit is 2 more than the last digit
a = e + 2
but less than the second digit
a < b
the second digit is the third multiple of 3
b = 9
while the fourth digit is the second multiple of 3
d = 6
the third digit is the difference between the second and the fourth digits
c = b - d
c = 9 - 6
c = 3
now we have: _ 9 3 6 _
Could be 89366, 79365, 69364............39361
what number